change in thermal energy (J)= mass (kg) x specific heat capacity (j/kg celcius) x change in temerature (celcius)
a 100 kg block of iron or an equal mass of water? the specific heat capacity of water is 4200 J/kg celcius and of iron is 450 J/kg celcius.
1
Expert's answer
2019-10-30T12:34:21-0400

Change in thermal energy for iron will be greater:


"mc_i \Delta T_i =mc_w \Delta T_w"

"\frac{\Delta T_i }{\Delta T_w }=\frac{c_w}{c_i}=\frac{4200}{450}=9.3>1"